Suppose you travel north for 65 kilometers then travel east 75 kilometers. How far are you from your starting point?
Answer by checkley75(3666)

THIS IS A RIGHT TRIANGLE PROBLEM WITH SIDES OF 65 & 75 KM.
NOW WE NEED TO FIND THE HYPOTENUSE WHICH IS A^2+B62=C^2 C BEING THE HYPOTENUSE.
65^2+75^2=C^2
4225+5625=C^2
9850=C^2
C=SQRT9850
C=99.25 KM. FOR THE HYPOTENUSE.